OpenStudy (anonymous):

Solve: 4(x + 1) = −4

OpenStudy (anonymous):

-1

OpenStudy (diyadiya):

Distribute: 4(x+1)= -4 4*x + 4*1 = -4 4x + 4 = -4

hero (hero):

@michelle1503, read the CoC

OpenStudy (anonymous):

sorry the what?

hero (hero):

4(x+1) = -4 4x + 4 = -4 4x = -4 - 4 4x = -8 x = -8/4 x = ?

hero (hero):

http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

-2

OpenStudy (diyadiya):

Right!!